3. Capital of 3 200 EUR was paid to the bank, the half-yearly interest rate is 2%, monthly capitalization. Calculate the amount you will get after 2 years
Answer #1

To calculate Feature value of 3200 monthly capitalised. At the rate of 2% for 2 years is.

FV= P(1+r/t)nxt

FV= 3200(1+0.02/12)whole power 2x12

FV= 3200( 1+0.0016)24

FV= 3200( 1.041)

FV= 3330